8. Given the following balanced reaction for the combustion of ethanol, an alternative fuel and gasoline additive C2H5OH(l) + 3O2(g) → 2CO2(g) + 3H2O(g) and the following table of standard enthalpies of formation. How much energy (in kJ) is released from the combustion of 1.25kg of ethanol?

(Continue on the next page)
Answer #1

Solution:

Fir the given reaction,

C2H5OH (l) + 3O2 (g) = 2CO2 (g) + 3H2O (g)

ΔH = Hf products - Hf reactants

= 2 x Hf CO2 + 3 x Hf H2O - 1 x Hf C2H5OH + 3 x Hf O2

= 2 x (- 393.5) + 3 x ( -241.8) - 1.x (-277.7) + 3 x O

= -787 - 725.4 + 277.5 = -1234.7 kJ / mol

Number of moles in 1.25 kg of ethanol

= mass / molar mass

= 1.25 kg / 46 g mol-1 = 1250 g / 46 g mol-1 = 27.174 mol

Thus,

Enthalpy change = 27.174 mol x -1234.7 kJ/mol

= -33551.7 kJ